OpenStudy (meilendurcer):

4/7-3/6? A. 1 B. 1/42 C. 1/14 D. 5/21

OpenStudy (anonymous):

\[ \frac{4}{7}*\frac{6}{6} - \frac{3}{6}*\frac{7}{7}\] \[ \frac{24}{42} - \frac{21}{42}=\frac{24-21}{42}\] \[ \frac{3}{42}=\frac{1}{14}\]